OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. Write the solution as an ordered pair (a, b). a - b = -5 and a + 2b = 1

OpenStudy (anonymous):

b=2 and a=-3

OpenStudy (campbell_st):

a - b = -5 (1) a + 2b = 1 (2) you need 2 x (1) + (2) 2a - 2b = -10 + a + 2b = 1 --------------- 3a = -9 a = -3 substitute into either equation to find b b = 2 hope it helps